# Breaking the Bottleneck with DiffuApriel: High-Throughput Diffusion LMs with Mamba Backbone

### v1

@@ 3.1 Preliminary
Here, $q_{0|t}(\boldsymbol{x}_{s}^{i}\mid\boldsymbol{x}_{t})$ denotes the model-provided distribution over the vocabulary for predicting a non-$\text{MASK}]$ token when $\boldsymbol{x}_{t}^{i}$ is masked. In conditional generation settings—e.g., generating a response $\boldsymbol{x}_{0}$ given a prompt pp—the reverse diffusion process in Equation 3 must be adapted. In this case, the model’s predictive distribution for unmasking a token becomes prompt-conditioned: $q_{0|t}(\boldsymbol{x}_{s}^{i}\mid\boldsymbol{x}_{t},p),$ reflecting that token predictions now depend on both the intermediate noised sequence and the conditioning prompt.
@@ 4.2 Inference Throughput and Modeling Quality
Figure 2: Inference throughput and Model’s Latency per forward pass vs. sequence length with a batch size of 1. It can be observed that inference throughput in DiffuApriel scales nearly linearly with $L$, whereas DiffuTran degrades super-linearly due to quadratic attention. At 1.3B scale, DiffuApriel and DiffuApriel+MLP yields 5.3x and 4.4x throughput improvement over DiffuTran respectively. Further hybrid architectures, DiffuApriel-H and DiffuApriel-H+MLP gives 2.8x and 2.6x throughput improvement over DiffuTran.
@@
We observe distinct scaling behaviors as the sequence length increases from L=32L{=}32 to L=65L{=}65K between the attention based diffusion model and our proposed DiffuApriel as shown in Figure 2. For DiffuApriel (we use Mamba-2 training kernel for both training and inference), inference throughput grows and saturates for longer sequences. This trend arises from their $\mathcal{O}(B\,L\,d)$ computational and memory complexity, where total latency $t(L)$ also scales linearly with $L$. Once the GPU reaches its peak memory bandwidth, the model becomes memory bound (Baruah et al., 2025), because the work (total tokens $\propto L$) and the latency ($t(L)\propto L$) scale equally, the resulting tokens-per-second rate ($T(L)\propto L/t(L)$) becomes approximately constant: $T_{\max}\approx M_{p}/\text{bytes per token}$, where $M_{p}$ is peak memory bandwidth per second (bytes/sec).
@@
In contrast, DiffuTran exhibit an initial rise followed by a sharp decline in throughput as sequence length increases, owing to their quadratic $\mathcal{O}(B\,L^{2}\,d)$ cost. The quadratic growth of the $QK^{\top}$ and softmax operations means the total latency $t(L)$ scales super-linearly ($t(L)\!\sim\!\beta BL^{2}$). Since throughput is defined as tokens per second ($T(L)\propto L/t(L)$), the $L$ growth in the numerator is overcome by the $L^{2}$ growth in the denominator, causing throughput to decay roughly as $T(L)\!\propto\!1/L$. This quadratic cost is further exacerbated by unfavorable GEMM (General Matrix–Matrix Multiplication) aspect ratios ($L\gg d_{h};\;d_{h}:\text{head dimension}$) at large $L$.
@@
For hybrid diffusion models interleaving state-space and attention layers, the overall complexity falls between the linear $\mathcal{O}(B\,L\,d)$ scaling of Mamba and the quadratic $\mathcal{O}(B\,L^{2}\,d)$ scaling of attention. With attention inserted every $K$ layers, the effective cost approximates $\mathcal{O}(B\,L\,d+B\,L^{2}\,d/K)$. Thus, the linear term dominates for moderate $L$, yielding near-linear throughput, while the quadratic component contributes mildly at larger $L$. This interleaving achieves a balance, preserving attention’s global context modeling while maintaining the high streaming efficiency of Mamba. These observations align with the theoretical cost model and confirm that Mamba-like architectures maintain stable inference efficiency under long-context regimes, whereas attention models face quadratic degradation beyond a critical sequence length.
@@ 5 Conclusion
This paper introduced a new architecture for masked DLMs by replacing Transformer denoisers with state space models. We presented DiffuApriel, to our knowledge the first diffusion LM built on a bidirectional Mamba backbone, and DiffuApriel-H, a hybrid variant that adds interleaved attention layers for complementary global context. Across 240M to 1.3B parameters, both models match or surpass the Transformer baseline DiffuTran while delivering significantly higher throughput and lower latency. Our results connect to a growing body of work exploring efficient linear-time mixers, including SSMs, DeltaNet-style token mixers, and hybrid architectures, underscoring that attention is not strictly required for iterative denoising. A remaining limitation of our approach is that performance advantages diminish on short contexts. Looking forward, masked diffusion with hybrid backbones can accommodate block-diffusion techniques in a complementary way. Exploring this integration is a compelling avenue for future work.

### v2

@@ 3.1 Preliminary
Here, $q_{0|t}(\boldsymbol{x}_{s}^{i}\mid\boldsymbol{x}_{t})$ denotes the model-provided distribution over the vocabulary for predicting a non-$\text{MASK}]$ token when $\boldsymbol{x}_{t}^{i}$ is masked. In conditional generation settings e.g., generating a response $\boldsymbol{x}_{0}$ given a prompt $p$, the reverse diffusion process in Equation 3 must be adapted. In this case, the model’s predictive distribution for unmasking a token becomes prompt-conditioned: $q_{0|t}(\boldsymbol{x}_{s}^{i}\mid\boldsymbol{x}_{t},p),$ reflecting that token predictions now depend on both the intermediate noised sequence and the conditioning prompt.
@@ 4.2 Inference Throughput and Modeling Quality
[…]
We observe distinct throughput behaviors as the sequence length increases from L=32L{=}32 to L=65L{=}65K, keeping the batch size as 1, between the attention-based diffusion model and our proposed DiffuApriel, as shown in Figure 2. We report end-to-end decoding throughput in tokens per second, computed as the number of generated tokens divided by the total wall-clock time from generation start to final token. We fix the decoding steps for DiffuApriel and DiffuApriel-H at 128, as in Nie et al. (2025), and set a block size of 32 with 2 denoising steps per block for DiffuTran+KV caching (Wu et al., 2025).
[…]
@@
For DiffuApriel (we use Mamba-2 training kernel for both training and inference), inference throughput grows and saturates for longer sequences. This trend arises from their $\mathcal{O}(B\,L\,d)$ computational and memory complexity, where total latency $t(L)$ also scales linearly with $L$. Once the GPU reaches its peak memory bandwidth, the model becomes memory bound (Baruah et al., 2025), because the work (total tokens $\propto L$) and the latency ($t(L)\propto L$) scale equally, the resulting tokens-per-second rate ($T(L)\propto L/t(L)$) becomes approximately constant: $T_{\max}\approx M_{p}/\text{bytes per token}$, where $M_{p}$ is peak memory bandwidth per second (bytes/sec).
@@
In contrast, DiffuTran exhibit an initial rise followed by a sharp decline in throughput as sequence length increases, owing to their quadratic $\mathcal{O}(B\,L^{2}\,d)$ cost. The quadratic growth of the $QK^{\top}$ and softmax operations means the total latency $t(L)$ scales super-linearly ($t(L)\!\sim\!\beta BL^{2}$). Since throughput is defined as tokens per second ($T(L)\propto L/t(L)$), the $L$ growth in the numerator is overcome by the $L^{2}$ growth in the denominator, causing throughput to decay roughly as $T(L)\!\propto\!1/L$. This quadratic cost is further exacerbated by unfavorable GEMM (General Matrix–Matrix Multiplication) aspect ratios ($L\gg d_{h};\;d_{h}:\text{head dimension}$) at large $L$. Furthermore, DiffuTran+KV caching (Wu et al., 2025) achieves higher throughput than DiffuApriel for shorter sequences (up to 2048 tokens), but degrades sharply beyond 10K tokens, consistent with Peng et al. (2025), eventually performing even worse than DiffuTran alone.
@@
For hybrid diffusion models interleaving state-space and attention layers, the overall complexity falls between the linear $\mathcal{O}(B\,L\,d)$ scaling of Mamba and the quadratic $\mathcal{O}(B\,L^{2}\,d)$ scaling of attention. With attention inserted every $K$ layers, the effective cost approximates $\mathcal{O}(B\,L\,d+B\,L^{2}\,d/K)$. Thus, the linear term dominates for moderate $L$, yielding near-linear throughput, while the quadratic component contributes mildly at larger $L$. This interleaving achieves a balance, preserving attention’s global context modeling while maintaining the high streaming efficiency of Mamba. These observations confirm that Mamba-like architectures maintain stable inference efficiency under long-context regimes, whereas attention models face quadratic degradation beyond a critical sequence length.
@@ 5 Conclusion
This paper introduced a new architecture for masked DLMs by replacing Transformer denoisers with state space models. We presented DiffuApriel, to our knowledge the first diffusion LM built on a bidirectional Mamba backbone, and DiffuApriel-H, a hybrid variant that adds interleaved attention layers for complementary global context. Across 240M to 1.3B parameters, both models match or surpass the Transformer baseline DiffuTran while delivering significantly higher throughput and lower latency. Our results connect to a growing body of work exploring efficient linear-time mixers, including SSMs, DeltaNet-style token mixers, and hybrid architectures, underscoring that attention is not strictly required for iterative denoising. A remaining limitation of our approach is that performance advantages diminish on short contexts. Looking ahead, investigating reasoning at scale within masked diffusion frameworks, particularly using Mamba and hybrid backbones, presents a promising and compelling direction for future research.